# Ironman: Accelerating Oblivious Transfer Extension for Privacy-Preserving AI with Near-Memory Processing

### v2

@@ 3.1. SPCOT Operation Analysis
PRG | Output size (bit) | Area () | Perf./Area Ratios | Power () | Power/Block Ratios | Security Bits |
AES-128 | 128 | 0.233 | 1 | 35.05 | 1 | 128 |
ChaCha8 | 512 | 0.215 | 4.491 | 45.34 | 3.092 | (Aumasson, 2019) |
(2) AES-based PRG is not optimal for customized protocol. In customized hardware scenarios, ChaCha-based PRGs exhibit better properties. As shown in Table 2, we evaluate different fully-pipelined PRGs from three perspectives: ① Area: ChaCha8 is a suitable choice for area efficiency, occupying , compared to for AES. ② Power: ChaCha8 consumes 45mW, compared to AES’s 32mW. However, ChaCha8 produces 4 blocks per cycle (512 bits) while AES produces only one block per cycle (128 bits). Therefore, ChaCha8 is more power-efficient per block. ③ Security: It is widely accepted that 128-bit security is sufficient for most cryptographic applications. According to (Aumasson, 2019), 7-round ChaCha provides around 248-bit security. These advantages have been well recognized in the literature (Lam et al., 2023; Xiaolin et al., 2025). However, many of these works overlook a key property of ChaCha-based PRGs: their ability to generate longer output lengths. This feature is particularly suited for m-ary tree expansion and can be leveraged to further reduce the number of PRG calls.
@@ 5.3. Index Sorting Algorithm for Memory-side Cache
To achieve further latency reduction, we introduce Row Look-ahead cooperated with Column Swapping. Figure 11(c) give a simple example of how Row Look-ahead works. In previous methods, elements were read from the sparse matrix row by row. Here, we introduce to overcome this limitation. This allows us to pre-fetch the indices of subsequent rows and access the corresponding vectors. If the elements indicated by these indices are already present in the current cache, we can access them in advance, thereby reducing the likelihood of a cache miss. Because the sparse matrix can remain fixed, we simulate memory-side cache behavior on the offline phase and store the sorted CSR format which have a better temporal and spatial locality.
@@ 6. Experiment
Category | Parameters |
Host Processor | 24 cores, 2.2GHz, 1.5MB L1 cache, |
48MB L2 cache, 71.5MB LLC cache | |
DRAM Module | DDR4-2400MHz 8GB, 64GB total size, |
4 Channels × 2 DIMMs × 2 Ranks, FR-FCFS | |
DRAM Timing Parameters | tRCD=16, tCL=16, tRP=16, tRC=55, |
tRRD_S=4, tRRD_L=6, tFAW=26, | |
@@ 6.1. Overall PCG-style OTE Generation Speedup
We implement the PCG-style OTE protocol on the GPU for a comprehensive comparison. By leveraging GPU acceleration,the OTE protocol achieves a latency of 108.8 ms and delivers a improvement in throughput. The latency breakdown shows that SPCOT and LPN account for and of the total latency, respectively. The observed differences in the breakdown compared to the CPU results, as shown in Figure 1(b), are primarily due to the significantly larger L1 and L2 caches in the A6000 GPU, which provide higher bandwidth for LPN operations. Compared to the GPU implementation, as shown in Figure 12, Ironman achieves a reduction in latency, along with an reduction in power consumption.
@@ 6.4. Operation-wise Performance Profiling
![[Uncaptioned image]](x14.png)
![[Uncaptioned image]](x15.png)
We demonstrate Ironman’s effectiveness in accelerating PPML applications by optimizing the OT protocol for nonlinear functions. Specifically, we benchmark Ironman on LayerNorm, GeLU, Softmax, and ReLU within two PPML frameworks: EzPC-SiRNN (Rathee et al., 2021b; Chandran et al., 2017) and Bolt (Pang et al., 2024a). As shown in Figure 15, Our experiments show the following key observations: (1) Significant Latency Reduction: Ironman achieves a to reduction in latency across different nonlinear functions and frameworks. (2) Framework-Agnostic Acceleration: Ironman achieves around a latency reduction across frameworks, primarily due to OT optimization.
@@ 6.6. Overall Area and Power Analysis
ChaCha8 Core | Ironman-NMP | Typical DRAM chip | ||
256KB Cache | 1MB Cache | |||
Area () | 0.215 | |||
Power () | 45.33 | 1.301 | 1.430 |
@@ Acknowledgements.
We sincerely thank the anonymous reviewers for their feedback. This work was supported in part by NSFC under Grant 62495102, Grant 92464104, and Grant 62125401, in part by the National Key Research and Development Program under Grant 2024YFB4505004, in part by Beijing Municipal Science and Technology Program under Grant Z241100004224015, and in part by 111 Project under Grant B18001.References

### v3

@@ 2.3.1. SPCOT
In addition to the AES encryption, requires the output of AES to be XORed with its input. However, as this final XOR step has a negligible computational overhead and can be integrated into the AES unit, we omit it in the subsequent discussion for simplicity.
@@ 3.1. SPCOT Operation Analysis
| PRG |
| AES-128 | 128 | 0.233 | 1 | 35.05 | 1 | ||||||||||
| ChaCha8 | 512 | 0.215 | 4.491 | 45.34 | 3.092 |
(2) AES-based PRG is not optimal for customized protocol. In customized hardware scenarios, ChaCha-based PRGs exhibit better properties. As shown in Table 2, we evaluate different fully-pipelined PRGs from three perspectives: ① Area: ChaCha8 is a suitable choice for area efficiency, occupying , compared to for AES. ② Power: ChaCha8 consumes 45mW, compared to AES’s 32mW. However, ChaCha8 produces 4 blocks per cycle (512 bits) while AES produces only one block per cycle (128 bits). Therefore, ChaCha8 is more power-efficient per block. ③ Security: The 128-bit security level is sufficient for most of real-word applications. According to (Aumasson, 2019), 7-round ChaCha provides around 248-bit security. These advantages have been well recognized in the literature (Lam et al., 2023; Xiaolin et al., 2025). However, many of these works overlook a key property of ChaCha-based PRGs: their ability to generate longer output lengths. This feature is particularly suited for m-ary tree expansion and can be leveraged to further reduce the number of PRG calls.
@@ 5.3. Index Sorting Algorithm for Memory-side Cache
To achieve further latency reduction, we introduce Row Look-ahead cooperated with Column Swapping. Figure 11(c) give a simple example of how Row Look-ahead works. In previous methods, elements were read from the sparse matrix row by row. Here, we introduce to overcome this limitation. This allows us to pre-fetch the indices of subsequent rows and access the corresponding elements. If the elements indicated by these indices are already present in the current cache, we can access them in advance, thereby reducing the likelihood of a cache miss. Because the sparse matrix can remain fixed, we simulate memory-side cache behavior on the offline phase and store the sorted CSR format which have a better temporal and spatial locality.
@@ 6. Experiment
| Category | Parameters | ||
| Host Processor | 24 cores, 2.2GHz, 1.5MB L1 cache, | ||
| 48MB L2 cache, 71.5MB LLC cache | |||
| DRAM Module | DDR4-2400MHz 8GB, 64GB total size, | ||
| 4 Channels × 2 DIMMs × 2 Ranks, FR-FCFS | |||
tRCD=16, tCL=16, tRP=16, tRC=55, | ||
| tRRD_S=4, tRRD_L=6, tFAW=26, | |||
@@ 6.1. Overall PCG-style OTE Generation Speedup
We implement the PCG-style OTE protocol on the GPU for a comprehensive comparison. By leveraging GPU acceleration,the OTE protocol delivers a improvement in throughput. The latency breakdown shows that SPCOT and LPN account for and of the total latency, respectively. The observed differences in the breakdown compared to the CPU results, as shown in Figure 1(b), are primarily due to the significantly larger L1 and L2 caches in the A6000 GPU, which provide higher bandwidth for LPN operations. Compared to the GPU implementation, as shown in Figure 12, Ironman achieves a reduction in latency, along with an reduction in power consumption.
@@ 6.4. Operation-wise Performance Profiling
We demonstrate Ironman’s effectiveness in accelerating PPML applications by optimizing the OT protocol for nonlinear functions. Specifically, we benchmark Ironman on LayerNorm, GeLU, Softmax, and ReLU within two PPML frameworks: EzPC-SiRNN (Rathee et al., 2021b; Chandran et al., 2017) and Bolt (Pang et al., 2024a). As shown in Figure 6.4, Our experiments show the following key observations: (1) Significant Latency Reduction: Ironman achieves a to reduction in latency across different nonlinear functions and frameworks. (2) Framework-Agnostic Acceleration: Ironman achieves around a latency reduction across frameworks, primarily due to OT optimization.
@@ 6.6. Overall Area and Power Analysis
Ironman-NMP |
| 256KB Cache | 1MB Cache | |||||||
| Area () | 0.215 | |||||||
| Power () | 45.33 | 1.301 | 1.430 | |||||
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (NSFC) under Grants 62495102, 92464104, and 62125401; the National Key Research and Development Program of China under Grant 2024YFB4505004; the Beijing Municipal Science and Technology Program under Grant Z241100004224015; the 111 Project under Grant B18001; and the Damo Academy Innovative Research Program.References